OpenStudy (anonymous):

your parents are considering renting yoh an apartment instead of paying room and board at your college. the month to month contract is $539.00 month and the annual contact is $467.00 month. if you break the annual contact there is a 2 month penalty

OpenStudy (texaschic101):

wait...what is the question ?

OpenStudy (anonymous):

your parents are considering renting yoh an apartment instead of paying room and board at your college. the month to month contract is $539.00 month and the annual contact is $467.00 month. if you break the annual contact there is a 2 month penalty

OpenStudy (texaschic101):

I still do not understand what you are looking for ?

OpenStudy (johnweldon1993):

Right, see we have the data...but no actual question was asked :)

OpenStudy (texaschic101):

month to month contract would cost you : 539 (12) = 6468/yr I do not understand...if you break the contract...you mean by leaving early ??

OpenStudy (anonymous):

How much do your parents save if you stay a whole year by entering into an annual Contact rather than month to month contract

OpenStudy (anonymous):

srry for the not understaning

OpenStudy (texaschic101):

If you break the annual contract....are you still paying for the whole year plus the 2 months ?

OpenStudy (anonymous):

Wait ? I'm confused

OpenStudy (texaschic101):

if so...then it would be 467 (12) + 467 (2) = 5604 + 934 = 6538/yr + penalty for the annual one

OpenStudy (texaschic101):

lets see what john says

OpenStudy (johnweldon1993):

Luckily we don't even need to worry about the penalty since the question states you will be staying the whole year...so it just boils down to \[\large \text{Month to Month} = 539\times 12 = ?\] \[\large \text{Annual} = 467 \times 12 = ?\] And then how much is saved from doing the annual rather than the month to month

OpenStudy (texaschic101):

tricky....they threw some info in there that wasn't even needed

OpenStudy (anonymous):

6468

OpenStudy (anonymous):

?

OpenStudy (anonymous):

John u still here

OpenStudy (texaschic101):

the 6468 is for the month to month contract

OpenStudy (anonymous):

ok thx u

OpenStudy (texaschic101):

the annual one is 467 x 12 then you take the difference and that is your answer

OpenStudy (johnweldon1993):

Yeah, sorry guys my openstudy began to lag big time! Hope everything made sense!